A spring of spring constant 50 N/m is compressed from its natural position through 1 cm. Find the work done by the spring-force on the agency compressing the spring.

Grade:12th pass

1 Answers

Apoorva Arora IIT Roorkee
9 years ago
The magnitude of the work is

\frac{1}{2}kx^{2} = \frac{1}{2}\times (50 N/m)\times (1cm)^{2} = (25 N/m)\times (1\times 10^{-2} m)^{2} = 2.5\times 10^{-3} J.

As tehe compressed spring will push the agency, the force will be opposite to the displacement of the point of application and the work will be negative. Thus, the work done by the spring-force is -2.5 mJ.